Find the first 3 terms, in ascending powers of π‘₯, of the binomial expansion of (2βˆ’π‘₯)4 and simplify each term.


1
Expert's answer
2022-06-02T08:54:21-0400
"(a+b)^4=\dbinom{4}{0}a^4+\dbinom{4}{1}a^3b+\dbinom{4}{2}a^2b^2""+\dbinom{4}{3}ab^3+\dbinom{4}{4}b^4""=a^4+4a^3b+6a^2b^2+4ab^3+b^4"





"(2-x)^4=16-32x+24x^2-8x^3+x^4"

"16, -32x, 24x^2"